OpenStudy (anonymous):

Among all pairs of numbers whose difference is 8, find a pair whose product is as small as possible. What is the minimum product?

OpenStudy (anonymous):

x-y=8 x*(-8+x) -8-x^2

myininaya (myininaya):

x-y=8 implies x=8+y P=xy=(8+y)y=8y+y^2 P'=8+2y 8+2y=0 implies y=-4 implies x=8-4=4

OpenStudy (anonymous):

other than zero... a-b = 8 a = 8+b ab = (8+b)b minimize b^2 + 8b 2b+8 = 0 2b = -8 b=-4

OpenStudy (anonymous):

Thank you all very much. I am new to this site so this helps a lot. I'll have more questions throughout the next 24 hours.